# Higgs-portal vector dark matter at a low reheating temperature

### v1

@@ I Introduction
As a matter of fact, in order for Big Bang Nucleosynthesis (BBN) to work Hannestad:2004px , it is sufficient to take the temperature of the SM bath particles as low as 4 MeV after the reheating process is complete. Recently, FIMP DM has been studied for the case with a low reheating temperature in Refs. Cosme:2023xpa ; Cosme:2024ndc ; Arcadi:2024wwg ; Boddy:2024vgt ; Arcadi:2024obp ; Lebedev:2024vor ; Lee:2024wes ; Belanger:2024yoj . In this case, compared to the standard FIMP DM, DM is much heavier than the reheating temperature, so the thermal production of the DM from the decays and scatterings of thermal particles would be Boltzmann-suppressed. Thus, we need sufficiently sizable couplings for FIMP DM, so there are new possibilities to test the FIMP DM models by direct detection experiments, indirect detection searches and collider searches. Simple scenarios of DM with s-wave annihilation channels have been studied, including Scalar DM, Fermion DM or Vector DM with the SM Higgs boson mediator Arcadi:2024wwg and Fermion DM with a very light or very heavy $Z^{\prime}$ mediator Boddy:2024vgt ; Arcadi:2024obp .
@@ 00 $1$
where $W_{D\,\mu\nu}=\partial_{\mu}W_{D\,\nu}-\partial_{\nu}W_{D\,\mu}$ and the covariant derivative in the above Lagrangian for the extra field is defined as $D_{\mu}=\partial_{\mu}-ig_{D}n_{A}W_{D\mu}$ where $g_{D}$, $W_{D}$, $n_{A}$ are gauge coupling, gauge boson and charge of the field associated with the additional gauge group $U(1)_{D}$. The potential takes the form,
@@ V Dark Matter Results
where $Y_{\alpha}=\frac{n_{\alpha}}{s(T)}$ is the co-moving number density of species $\alpha$. The decay and scattering contributions to the total amount of DM density have been displayed in Eq. 21, $M_{sc}$ is any mass scale that can be SM or BSM Higgs mass, $z=\frac{M_{sc}}{T}$, $g_{*}$ is the effective relativistic d.o.f of the Universe, and the local entropy is given by
@@
where the vertex factor $g_{h_{i}W_{D}W_{D}}$ is defined earlier in Eq. 20.
@@
In the left panel (LP) and right panel (RP) of Fig. 4, we have shown the variation in $\left(\frac{n_{eq}\langle\sigma v\rangle_{W_{D}W_{D}}}{H},g_{D}\right)$ and $\left(\frac{n_{eq}\langle\sigma v\rangle_{W_{D}W_{D}}}{H},T_{R}\right)$ planes for three different values of $T_{R}/M_{W_{D}}$ and $g_{D}$, respectively. The quantity $n_{eq}\langle\sigma v\rangle_{W_{D}W_{D}}/H$ measures the DM chemical equilibrium, depending on whether it is greater than $1$ or less than $1$. In the LP, we have shown the variation with three values of $T_{R}/M_{W_{D}}$ and we notice that as it decreases, DM remains in chemical equilibrium for higher values of the gauge coupling (green to blue line). This simply happens because the number density $n_{eq}\propto e^{-\frac{M_{W_{D}}}{T_{R}}}$ and the cross-section times velocity $\langle\sigma v\rangle_{W_{D}W_{D}}\propto g^{2}_{D}e^{-\frac{2M_{W_{D}}}{T_{R}}}$; therefore, the reduction in $n_{eq}$ and $\langle\sigma v\rangle_{W_{D}W_{D}}$ is compensated by the increase in $\langle\sigma v\rangle$ for different values of $T_{R}/M_{W_{D}}$ to reach chemical equilibrium. In the RP, we have shown the variation with $T_{R}$ for three different values of the gauge coupling $g_{D}$ while the DM mass is kept at $M_{W_{D}}=100$ GeV for both plots. We can see that for lower values of $g_{D}$, higher values of $T_{R}$ are needed to reach chemical equilibrium, which is equivalent to an increase in the number density of DM. Moreover, as the reheating temperature $T_{R}$ approaches the DM mass, i.e., $M_{W_{D}}=100$ GeV, we observe a flattening of $n_{eq}\langle\sigma v\rangle$ which implies less exponential suppression in the number density and $\langle\sigma v\rangle$. Both plots indicate that increasing $g_{D}$ allows DM to reach chemical equilibrium easily but can be avoided by choosing the lower values of $T_{R}/M_{W_{D}}$. This property of associating a low reheating temperature with higher values of the gauge coupling, while avoiding DM chemical equilibrium, allows for detection prospects of freeze-in DM.
@@
In Fig. 5, we have shown the variation of DM relic density with $g_{D}$ and $T_{R}$ for three different values of $T_{R}/M_{W_{D}}$ and $g_{D}$ in the LP and RP, respectively. In both plots, we have kept the DM mass fixed at $M_{W_{D}}=100$ GeV as mentioned in the caption of Fig. 4. In the LP, we can see a clear distinction between DM production via the freeze-in and freeze-out mechanisms. In the case of the freeze-in mechanism, the DM production rate increases with the gauge coupling value, whereas for the freeze-out mechanism, we expect the opposite behaviour. To obtain the plot, we checked whether DM reaches chemical equilibrium using our own developed code in micrOMEGAs. After that, we solved the Boltzmann equation accordingly using micrOMEGAs. Notice that we can see the change in the slopes at a certain point. This comes from the change of DM production from freeze-in to freeze-out depending on $g_{D}$ and the $T_{R}/M_{W_{D}}$ values. The important point to take from the plot is that for lower values of $T_{R}$, we can obtain the correct DM density for larger values of $g_{D}$. Hence, there are detection prospects for DM even if it is produced via the freeze-in mechanism. In the RP, we have shown the variation of DM density with the reheating temperature $T_{R}$. We can see that for different values of $g_{D}$, the DM relic density varies differently. The three values of $g_{D}=1,10^{-3}$ and $10^{-5}$, DM reaches thermal equilibrium at different values of $T_{R}$ and after that the DM relic density freezes out, as seen from the flat line. Once the DM reaches thermal equilibrium its density does not change because it is independent of $T_{R}$ contrary to the freeze-in mechanism. The density by freeze out mechanism changes when the interaction strength change which is fixed for each line in the RP and we have obtained three flat lines for three different values of $g_{D}$.
@@
In Fig. 9, we present scattered plots in the $(M_{W_{D}},\leavevmode\nobreak\ \sigma_{\rm SI})$ plane, where the colour variation in the LP corresponds to different values of the gauge coupling $g_{D}$, while in the RP, it represents the reheating temperature $T_{R}$. The analytical expression for the SIDD cross section for the VDM is given in Eq. 17. In the LP, we see from Eq. (17) that the SIDD cross section is proportional to the square of the gauge coupling, which is evident from the colour variation. An interesting aspect of the SIDD cross-section at low reheating temperatures is that only a small portion of the parameter space has been explored by LUX-ZEPLIN, leaving a large region open for future exploration—this stands in sharp contrast to present-day WIMP DM searches. In the RP, we observe that the SIDD cross section is independent of the reheating temperature, as indicated by Eq. (17). The vertical variation of $T_{R}$ with DM mass arises solely from relic density constraints. As discussed in the LP of Fig. 6, we have considered points where the relic density falls within the range $10^{-3}\leq\Omega_{DM}h^{2}\leq 0.12$ to obtain more data points within a reasonable time frame. By slightly adjusting the reheating temperature, we can achieve the total DM density for any value in the scattered plots. Therefore, we have not multiplied by the fraction of DM and have instead considered the bound equivalent to 100% DM. In both the plots, we have shown the recent bound from LUX-ZEPLIN LZ:2024zvo (green solid line) and future projection from DARWIN experiment DARWIN:2016hyl (blue dashed line) and $\nu-$fog (cyan dashed-dot line).
@@
In Fig. 10, we show the indirect detection bounds for DM. In the LP, we present results for the $b\bar{b}$ channel, while in the RP, we show results for the $WW$ channel. The analytical expressions for VDM annihilation to $b\bar{b}$ and $WW$ channels are given in Eq. 19. The colour variation in the LP represents different values of the reheating temperature $T_{R}$, whereas in the RP, it corresponds to different values of the gauge coupling $g_{D}$. In the LP, we observe that the indirect detection cross-section does not depend on the reheating temperature, as it only affects the DM relic density, leading to a vertical correlation. In the RP, the colour variation highlights different values of $g_{D}$, and we see that the cross-section increases with the gauge coupling. In both plots, we include future prospects for the CTA and Fermi-dSphs data, which are close to probing the present model parameter space. We anticipate that future, more sensitive indirect detection experiments or upgraded versions of the proposed experiments will further explore this model.

### v2

@@ I Introduction
As a matter of fact, in order for Big Bang Nucleosynthesis (BBN) to work Hannestad:2004px , it is sufficient to take the temperature of the SM bath particles as low as 4 MeV after the reheating process is complete. Recently, FIMP DM has been studied for the case with a low reheating temperature in Refs. Bhattiprolu:2022sdd ; Cosme:2023xpa ; Cosme:2024ndc ; Arcadi:2024wwg ; Boddy:2024vgt ; Arcadi:2024obp ; Lebedev:2024vor ; Lee:2024wes ; Belanger:2024yoj . In this case, compared to the standard FIMP DM, DM is much heavier than the reheating temperature, so the thermal production of the DM from the decays and scatterings of thermal particles would be Boltzmann-suppressed. Thus, we need sufficiently sizable couplings for FIMP DM, so there are new possibilities to test the FIMP DM models by direct detection experiments, indirect detection searches and collider searches. Simple scenarios of DM with s-wave annihilation channels have been studied, including Scalar DM, Fermion DM or Vector DM with the SM Higgs boson mediator Arcadi:2024wwg and Fermion DM with a very light or very heavy $Z^{\prime}$ mediator Boddy:2024vgt ; Arcadi:2024obp .
@@ 00 $1$
where $W_{D\,\mu\nu}=\partial_{\mu}W_{D\,\nu}-\partial_{\nu}W_{D\,\mu}$ and the covariant derivative in the above Lagrangian for the extra field is defined as $D_{\mu}=\partial_{\mu}-ig_{D}n_{A}W_{D\mu}$ where $g_{D}$, $W_{D}$, $n_{A}$ are gauge coupling, gauge boson and charge of the field associated with the additional gauge group $U(1)_{D}$ 111 For $SU(2)_{D}$ dark matter and connection to $U(1)$ VDM, see Ref. Hambye:2008bq . . The potential takes the form,
@@ V Dark Matter Results
where $Y_{\alpha}=\frac{n_{\alpha}}{s(T)}$ is the co-moving number density of species $\alpha$. The decay and scattering contributions to the total amount of DM density have been displayed in Eq. (21), $M_{sc}$ is any mass scale that can be SM or BSM Higgs mass, $z=\frac{M_{sc}}{T}$, $g_{*}$ is the effective relativistic d.o.f of the Universe, and the local entropy is given by
@@
where the vertex factor $g_{h_{i}W_{D}W_{D}}$ is defined earlier in Eq. (20).
@@
In the left panel (LP) and right panel (RP) of Fig. 4, we have shown the variation in $\left(\frac{n_{eq}\langle\sigma v\rangle_{W_{D}W_{D}}}{H},g_{D}\right)$ and $\left(\frac{n_{eq}\langle\sigma v\rangle_{W_{D}W_{D}}}{H},T_{R}\right)$ planes for three different values of $T_{R}/M_{W_{D}}$ and $g_{D}$, respectively. The quantity $n_{eq}\langle\sigma v\rangle_{W_{D}W_{D}}/H$ measures the DM chemical equilibrium, depending on whether it is greater than $1$ or less than $1$. In the LP, we have shown the variation with three values of $T_{R}/M_{W_{D}}$ and we notice that as it decreases, DM remains in chemical equilibrium for higher values of the gauge coupling (green to blue line). This simply happens because the number density $n_{eq}\propto e^{-\frac{M_{W_{D}}}{T_{R}}}$ and the cross-section times velocity $\langle\sigma v\rangle_{W_{D}W_{D}}\propto g^{2}_{D}e^{-\frac{2M_{W_{D}}}{T_{R}}}$; therefore, the reduction in $n_{eq}$ and $\langle\sigma v\rangle_{W_{D}W_{D}}$ is compensated by the increase in $\langle\sigma v\rangle$ for different values of $T_{R}/M_{W_{D}}$ to reach chemical equilibrium. Furthermore, when $g_{D}>0.1$, a plateau appears at $T_{R}/M_{W_{D}}=$1 and 0.5, which results from the mutual cancellation between the Higgs-like resonance regions of $h_{1}$ and $h_{2}$. In the RP, we have shown the variation with $T_{R}$ for three different values of the gauge coupling $g_{D}$ while the DM mass is kept at $M_{W_{D}}=100$ GeV for both plots. We can see that for lower values of $g_{D}$, higher values of $T_{R}$ are needed to reach chemical equilibrium, which is equivalent to an increase in the number density of DM. Moreover, as the reheating temperature $T_{R}$ approaches the DM mass, i.e., $M_{W_{D}}=100$ GeV, we observe a flattening of $n_{eq}\langle\sigma v\rangle$ which implies less exponential suppression in the number density and $\langle\sigma v\rangle$. Both plots indicate that increasing $g_{D}$ allows DM to reach chemical equilibrium easily but can be avoided by choosing the lower values of $T_{R}/M_{W_{D}}$. This property of associating a low reheating temperature with higher values of the gauge coupling, while avoiding DM chemical equilibrium, allows for detection prospects of freeze-in DM. Additionally, a kink appears around $T_{R}\sim 20$ GeV, corresponding to the transition of dominant DM annihilation channels from $WW,\leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ ZZ$ (below) to $WW,\leavevmode\nobreak\ h_{1}h_{1},\leavevmode\nobreak\ ZZ$, and $t\bar{t}$ (above).
@@
In Fig. 5, we have shown the variation of DM relic density with $g_{D}$ and $T_{R}$ for three different values of $T_{R}/M_{W_{D}}$ and $g_{D}$ in the LP and RP, respectively. In both plots, we have kept the DM mass fixed at $M_{W_{D}}=100$ GeV as mentioned in the caption of Fig. 4. In the LP, we can see a clear distinction between DM production via the freeze-in and freeze-out mechanisms. In the case of the freeze-in mechanism, the DM production rate increases with the gauge coupling value, whereas for the freeze-out mechanism, we expect the opposite behaviour. To obtain the plot, we checked whether DM reaches chemical equilibrium using our own developed code in micrOMEGAs. After that, we solved the Boltzmann equation accordingly using micrOMEGAs. Notice that we can see the change in the slopes at a certain point. This comes from the change of DM production from freeze-in to freeze-out depending on $g_{D}$ and the $T_{R}/M_{W_{D}}$ values. The important point to take from the plot is that for lower values of $T_{R}$, we can obtain the correct DM density for larger values of $g_{D}$. Hence, there are detection prospects for DM even if it is produced via the freeze-in mechanism. In the RP, we have shown the variation of DM density with the reheating temperature $T_{R}$. We can see that for different values of $g_{D}$, the DM relic density varies differently. The three values of $g_{D}=1,10^{-3}$ and $10^{-5}$, DM reaches thermal equilibrium at different values of $T_{R}$ and after that the DM relic density freezes out, as seen from the flat line. Once the DM reaches thermal equilibrium its density does not change because it is independent of $T_{R}$ contrary to the freeze-in mechanism. The density by freeze out mechanism changes when the interaction strength change which is fixed for each line in the RP and we have obtained three flat lines for three different values of $g_{D}$. In both plots, we have first computed the DM relic density using the freeze-out mechanism until its interaction strength is greater than the Hubble rate. We then produced the DM density using the freeze-in mechanism and extended it until it intersects with the freeze-out regime. It is to be noted that around the intersection point between WIMP and FIMP regimes, we need to calculate the temperature of the dark sector and compare it with the SM bath temperature. Based on the temperature comparison, we have to use the freeze-in or freeze-out mechanism for DM production, which we left for the future work. Finally, our study has been carried out when DM is completely out of equilibrium, and the intersection points between the WIMP and FIMP regimes do not affect our conclusions.
@@
In Fig. 9, we present scattered plots in the $(M_{W_{D}},\leavevmode\nobreak\ \sigma_{\rm SI})$ plane, where the colour variation in the LP corresponds to different values of the gauge coupling $g_{D}$, while in the RP, it represents the reheating temperature $T_{R}$. The analytical expression for the SIDD cross section for the VDM is given in Eq. (17). In the LP, we see from Eq. (17) that the SIDD cross section is proportional to the square of the gauge coupling, which is evident from the colour variation. An interesting aspect of the SIDD cross-section at low reheating temperatures is that only a small portion of the parameter space has been explored by LUX-ZEPLIN, leaving a large region open for future exploration—this stands in sharp contrast to present-day WIMP DM searches. In the RP, we observe that the SIDD cross section is independent of the reheating temperature, as indicated by Eq. (17). The vertical variation of $T_{R}$ with DM mass arises solely from relic density constraints. As discussed in the LP of Fig. 6, we have considered points where the relic density falls within the range $10^{-3}\leq\Omega_{DM}h^{2}\leq 0.12$ to obtain more data points within a reasonable time frame. By slightly adjusting the reheating temperature, we can achieve the total DM density for any value in the scattered plots. Therefore, we have not multiplied by the fraction of DM and have instead considered the bound equivalent to 100% DM. In both the plots, we have shown the recent bound from LUX-ZEPLIN LZ:2024zvo (green solid line) and future projection from DARWIN experiment DARWIN:2016hyl (blue dashed line) and $\nu-$fog (cyan dashed-dot line).
@@
In Fig. 10, we show the indirect detection bounds for DM. In the LP, we present results for the $b\bar{b}$ channel, while in the RP, we show results for the $WW$ channel. The analytical expressions for VDM annihilation to $b\bar{b}$ and $WW$ channels are given in Eq. (19). The colour variation in the LP represents different values of the reheating temperature $T_{R}$, whereas in the RP, it corresponds to different values of the gauge coupling $g_{D}$. In the LP, we observe that the indirect detection cross-section does not depend on the reheating temperature, as it only affects the DM relic density, leading to a vertical correlation. In the RP, the colour variation highlights different values of $g_{D}$, and we see that the cross-section increases with the gauge coupling. In both plots, we include future prospects for the CTA and Fermi-dSphs data, which are close to probing the present model parameter space. We anticipate that future, more sensitive indirect detection experiments or upgraded versions of the proposed experiments will further explore this model.